OpenStudy (anonymous):

how would i begin to solve this problem? 1/2+4m=3m-5/2

jimthompson5910 (jim_thompson5910):

I'm guessing the equation is \[\Large \frac{1}{2}+4m=3m-\frac{5}{2}\] right?

OpenStudy (anonymous):

yes what do i do first?

jimthompson5910 (jim_thompson5910):

You want to get rid of the fractions first

jimthompson5910 (jim_thompson5910):

So multiply every term by the LCD 2

OpenStudy (anonymous):

like 1/2(4m*2)(3*2)(-5*2)

jimthompson5910 (jim_thompson5910):

you're on the right track Multiply 1/2 by 2 to get 1 Multiply 4m by 2 to get 8m Multiply 3m by 2 to get 6m Multiply -5/2 by 2 to get -5 So after doing this, the equation becomes 1 + 8m = 6m - 5

jimthompson5910 (jim_thompson5910):

Do you see how I got that?

OpenStudy (anonymous):

not that great in math so im trying to study your work give me a min please.

jimthompson5910 (jim_thompson5910):

sure thing, take all the time you need

OpenStudy (anonymous):

so next i would have (8m+1)+(-1-6m) =(6m-5)+(-1-6m) or something like that

jimthompson5910 (jim_thompson5910):

yes, you are subtracting 1 from both sides and 6m from both sides to get 8m - 6m = -5 - 1

OpenStudy (anonymous):

im sorry im lost now

jimthompson5910 (jim_thompson5910):

how did you get (8m+1)+(-1-6m) =(6m-5)+(-1-6m) ?